OpenStudy (anonymous):

A store is having a 10% off sale. It gives an additional 20% off for an item that is above $100 after the first 10% discount.A jacket costs $124.99 if you buy the jacket, how much will you save in total? *SHOW WORK*

OpenStudy (lilg132):

124.99 / 100 = 1.2499 (This is 1% of the price) 1.2499 * 10 = 12.499 = 12.50 (This is 10% of the price) so you minus the price with 10% 124.99 - 12.50 = 112.49 is the price after 10% discount now to get the further 20% off the price you will use the same method to work out 20% of the price and then minus the price with 20%. 112.49 / 100 = 1.1249 (This is 1% of the price) 1.1249 * 20 = 22.498 = 22.50 (This is 20% of the price) You minus the price with 20% so you will get 112.49 - 22.50 = $89.99 is the total price after all discounts so the total amount you would have saved would be worked out by getting the original price and minus it by the price after all discounts so you will get 124.99 - 89.99 = $35 is the total amount you will save hope this helps you understand dont forget to click good answer :)